At the time it defaulted on its interest payments and filed for bankruptcy, Hanson Hardware Stores, Inc. had the balance sheet shown in the Exam Two Excel file (in thousands of dollars). The court, after trying unsuccessfully to reorganize the firm, decided that the only recourse was liquidation under Chapter 7. Sale of the fixed assets, which were pledged as collateral to the mortgage bondholders, were expected to bring in 75% of their book value. the current assets were expected to sell for 50% of their book value. Trustee’s costs totaled $40,000; no single worker was due more than $2,000 in wages and there were no unfunded pension plan liabilities.

              a. After the assets are liquidated, what is the total dollar amount available for distribution to all claimants?

              b. What is the total of all creditor and trustee claims?

              c. Will the preferred and common stockholders receive any distributions from the liquidation of the assets? Why or why not?

Hanson Hardware Stores, Inc,
Balance Sheet As of December 31, 2017
(Dollars in Thousands)
-------------2017------------
DOLLARS % OF ASSETS
ASSETS:
Current Assets 700 42.42%
Net Fixed Assets 950 57.58%
    TOTAL ASSETS 1,650 100.00%
LIABILITIES AND EQUITY:
Accounts Payable 365 22.12%
Accrued Taxes 30 1.82%
Accrued Wages 95 5.76%
Notes Payable 220 13.33%
    Total Current Liabilities 710 43.03%
Mortgage Bonds 800 48.48%
Subordinated Debenture Bonds 200 12.12%
    TOTAL LIABILITIES 1,710 103.64%
Common Stock 50 3.03%
Retained Earnings -110 -6.67%
    TOTAL STOCKHOLDERS' EQUITY -60 -3.64%
    TOTAL LIABILITIES AND EQUITY 1,650 100.00%
Answer #1

a). Total dollar amount available after liquidation = 50%*current assets book value + 75%*fixed assets book value

= 50%*700,000 + 75%*950,000 = $1,062,500

b).Total of trustee and all creditors' claims = total liabilities + trustee costs = 1,710,000 + 40,000 = $1,750,000

c). Preferred and common shareholders will not received anything as the entire amount after liquidation is less than the total claims and so, will be spent on settling those first. There will not be anything left for the shareholders.